### v1

Abstract
We propose the Regge trajectory relations for the fully heavy pentaquark $P_{cc\bar{c}bb}$ utilizing both diquark and triquark Regge trajectory relations. Using these new relations, we discuss four series of Regge trajectories: the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-trajectories. We provide rough estimates for the masses of the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-excited states. Except for the $\lambda_{1}$-trajectories, the complete forms of the other three series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ are lengthy and cumbersome. We show that the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can not be obtained by simply imitating the meson Regge trajectories because mesons have no substructures. To derive these trajectories, pentaquark’s structure and substructure should be taken into consideration. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will remain unclear, and the predictive power of the Regge trajectories would be compromised. Moreover, we show that the lengthy complete forms of the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can be well approximated by the simple fitted formulas. Four series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ all exhibit a behavior of $M{\sim}x^{2/3}$, where $x=n_{r_{1}},n_{r_{2}},l_{1},l_{2},N_{r_{1}},N_{r_{2}},L_{1},L_{2}$. All four series of trajectories exhibit concave downward behavior in the $(M^{2},\,x)$ plane.
$\lambda$-trajectory, $\rho$-trajectory, pentaquark, mass
I Introduction
As a type of exotic hadrons, pentaquarks will enhance the study of hadrons and will provide new probes for understanding QCD LHCb:2015yax ; ParticleDataGroup:2024cfk ; Jaffe:2004ph ; Ali:2017jda ; Liu:2019zoy ; Karliner:2003dt ; Roca:2015dva ; Lebed:2015tna ; Okiharu:2004wy ; Lebed:2016hpi ; Olsen:2017bmm ; Brambilla:2019esw . In 2015, the first observation of pentaquark was reported by the LHCb collaboration in $\Lambda^{0}_{b}\to J/{\psi}K^{-}p$ decays LHCb:2015yax . In recent years, exotica states containing fully heavy quarks have been observed. In 2020, a narrow structure around 6.9 ${\rm{GeV}/c^{2}}$ matching the lineshape of a resonance and a broad structure just above twice the $J/\psi$ mass were observed by the LHCb Collaboration LHCb:2020bwg . In 2023, the ATLAS Collaboration observed an excess of di-charmonium events in the four-muon final state with the ATLAS detector ATLAS:2023bft . In 2023, the CMS Collaboration reported new structures in the $J/{\psi}J/\psi$ mass spectrum in Proton-Proton Collisions at $\sqrt{s}=13\,{\rm TeV}$ CMS:2023owd .
These observations suggest the existence of fully heavy pentaquarks and motivate the studies on the fully heavy pentaquarks. Theoretical approaches include the quark model Gordillo:2024blx ; Liang:2024met ; Gordillo:2023tnz , chromomagnetic interaction model An:2020jix ; An:2022fvs , extended Gursey-Radicati formalism Sharma:2025adr , quark delocalization color screening model Yan:2021glh , MIT bag model Zhang:2023hmg , sum rules Wang:2021xao ; Azizi:2025fmx ; Zhang:2020vpz , lattice-QCD inspired quark model Yang:2022bfu , and the effective quark mass and screened charge scheme Rashmi:2024ako , among others.
The Regge trajectory [^1] A Regge trajectory of bound states is generally expressed as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$ Chen:2022flh ; Chen:2021kfw , where $M$ is the mass of the bound state, $l$ is the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, plots in the $(M,\,x)$ plane Xie:2024lfo , $(M-m_{R},\,x)$ plane Chen:2023cws , $(M,\,(x+c_{0})^{\nu})$ plane Burns:2010qq , $(M^{2},\,x)$ plane Chen:2018nnr , $((M-m_{R})^{2},\,x)$ plane Chen:2023djq ; Chen:2023web or $((M-m_{R})^{1/{\nu}},\,x)$ plane Xie:2024dfe , are all commonly referred as Chew-Frautschi plots. Regge trajectories can be plotted in these various planes. is one of the effective approaches widely used in the study of hadron spectra Burns:2010qq ; Regge:1959mz ; Chew:1962eu ; Nambu:1978bd ; Gross:2022hyw ; Brodsky:2006uq ; Nielsen:2018uyn ; brau:04bs ; Brisudova:1999ut ; Guo:2008he ; Ebert:2009ub ; Irving:1977ea ; Collins:1971ff ; Inopin:1999nf ; Afonin:2014nya ; MartinContreras:2020cyg ; Sergeenko:1994ck ; Veseli:1996gy ; Inopin:2001ub ; Wilczek:2004im ; Sonnenschein:2018fph ; MartinContreras:2023oqs ; Roper:2024ovj ; G:2024zkc ; Oudichhya:2024ikt . Few studies have applied the Regge trajectory approach to the discussion of pentaquarks. In Ref. Ghosh:2017cck , Regge trajectories for the pentaquarks containing one heavy quark were investigated. In Ref. Sindhu:2023oqo , Regge trajectories for the pentaquarks containing two charm quarks were discussed. In both of these references, only one series of Regge trajectories are discussed. To our knowledge, there have been no systematic studies up to now addressing both two series of $\lambda$-trajectories and two series of $\rho$-trajectories for pentaquarks. In the diquark-triquark picture, a pentaquark consists of a diquark and a triquark, where the triquark is composed of an antiquark and a diquark (see Fig. 1). By employing the diquark Regge trajectory relations Feng:2023txx and the triquark Regge trajectory relations Song:2024bkj , we attempt to apply the Regge trajectory approach to the pentaquark $P_{cc\bar{c}bb}$. The masses of two $\lambda$-mode and two $\rho$-mode excited states are roughly estimated. Four series of Regge trajectories–the $\lambda_{1}$-, $\lambda_{2}$-, $\rho_{1}$-, and $\rho_{2}$-trajectories–are investigated.
The paper is organized as follows: In Sec. II, the Regge trajectory relations for the the fully heavy pentaquarks are proposed. In Sec. III, four series of masses and four series of Regge trajectories are presented. In Sec. IV, the discussions are given. The conclusions are presented in Sec. V.
II Regge trajectory relations for the fully heavy pentaquarks
In this section, by utilizing the diquark Regge trajectories Feng:2023txx and the triquark Regge trajectories Song:2024bkj , we propose the pentaquark Regge trajectories, which can be employed to discuss both $\lambda$- and $\rho$-trajectories.
II.1 Preliminary
Figure 1: Schematic diagram of a pentaquark in the diquark-triquark picture. The left grey part represents the diquark $1$, composed of two quarks ($q_{1}$ and $q_{2}$). The right grey part represents a triquark composed of one antiquark ($\bar{q}_{3}$) and diquark $2$. Diquark $2$ is composed of quarks $q_{4}$ and $q_{5}$. The circles denote the quarks and the antiquark.
In the diquark-triquark picture, a pentaquark consists of one diquark and one triquark (see Fig. 1). $\rho_{1}$ and $\rho_{2}$ separates the quarks in the diquark $1$ and $2$, respectively. $\lambda_{1}$ corresponds to the separation between diquark $1$ and the triquark while $\lambda_{2}$ corresponds to the separation between the antiquark $\bar{q}_{3}$ and diquark $2$. There exist four excited modes: the $\rho_{1}$-mode involves the radial and orbital excitation in the diquark $1$; the $\rho_{2}$-mode involves excitation in the diquark $2$; the $\lambda_{1}$-mode involves the radial or orbital excitation between the diquark $1$ and triquark; and the $\lambda_{2}$-mode involves excitation between the antiquark and diquark $2$. Consequently, there are four series of Regge trajectories: two series of $\rho$-Regge trajectories and two series of $\lambda$-Regge trajectories.
In $SU_{c}(3)$, $3_{c}\otimes 3_{c}=\bar{3}_{c}\oplus{6}_{c}$. There is attraction between quark pairs $(q_{1}q_{2})$ in the color antitriplet channel. Only the $\bar{3}_{c}$ representation of $SU_{c}(3)$ is considered in the present work; the $6_{c}$ representation is not considered. Similarly, in the diquark picture, a triquark $(\bar{q}_{3}(q_{4}q_{5}))$ is regarded as a bound state consisting of a diquark $(q_{4}q_{5})$ in $\bar{3}_{c}$ and an antiquark $\bar{q}_{3}$ also in $\bar{3}_{c}$ Lebed:2015tna . In the triquark-diquark model, the pentaquarks under consideration are composed of a diquark and a triquark in the color $\bar{3}_{c}$ and $3_{c}$ configurations.
In the diquark-triquark picture, the state of the pentaquark is denoted as
$\left((q_{1}q_{2})^{{\bar{3}_{c}}}_{n_{1}^{2s_{1}+1}l_{1j_{1}}}\left(\bar{q}_{3}(q_{4}q_{5})^{{\bar{3}_{c}}}_{n_{2}^{2s_{2}+1}l_{2j_{2}}}\right)^{{{3}_{c}}}_{N_{2}^{2s_{3}+1}L_{2J_{2}}}\right)^{1_{c}}_{N_{1}^{2s_{4}+1}L_{1J_{1}}},$ (1)
[…]
In the diquark-triquark picture, the fully heavy pentaquark $P_{cc\bar{c}bb}$ has three configurations: $(cc)(\bar{c}(bb))$, $(bb)(\bar{c}(cc))$, and $(bc)(\bar{c}(bc))$. In the first configuration, $P_{cc\bar{c}bb}$ consists of a diquark $(cc)$ and a triquark $(\bar{c}(bb))$, where the triquark is composed of a antiquark $\bar{c}$ and a diquark $(bb)$. In the second configuration, $P_{cc\bar{c}bb}$ is composed of a diquark $(bb)$ and a triquark $(\bar{c}(cc))$. In the third configuration, $P_{cc\bar{c}bb}$ consists two clusters: $(bc)$ and $(\bar{c}(bc))$. Due to mode mixing in the $(bc)(\bar{c}(bc))$ configuration, the corresponding Regge trajectories are complex. This configuration is not considered in the present work.
II.2 Regge trajectory relations for the fully heavy pentaquarks
The spinless Salpeter equation Godfrey:1985xj ; Capstick:1986ter ; Ferretti:2019zyh ; Bedolla:2019zwg ; Durand:1981my ; Durand:1983bg ; Lichtenberg:1982jp ; Jacobs:1986gv is given by
$M\Psi_{d,t,p}({{\bf r}})=\left(\omega_{1}+\omega_{2}\right)\Psi_{d,t,p}({{\bf r}})+V_{d,t,p}\Psi_{d,t,p}({{\bf r}}),$ (2)
where $M$ is the bound state mass (diquark, triquark or pentaquark). $\Psi_{d,t,p}({{\bf r}})$ denotes the wave function of the diquark, triquark, and pentaquark, respectively. $V_{d,t,p}$ represents the diquark, triquark, and pentaquark potentials, respectively (see Eq. (4)). $\omega_{1}$ is the relativistic energy of constituent $1$ (quark, diquark, or antiquark), and $\omega_{2}$ is of constituent $2$ (quark, diquark, or triquark),
$\omega_{i}=\sqrt{m_{i}^{2}+{\bf p}^{2}}=\sqrt{m_{i}^{2}-\Delta}\;\;(i=1,2).$ (3)
Here, $m_{1}$ and $m_{2}$ are the effective masses of constituent $1$ and $2$, respectively.
The effect of the finite size of the diquark is treated differently. In Ref. Faustov:2021hjs , the size of diquark is taken into account through appropriate form factors, while sometimes the diquark is treated as pointlike Ferretti:2019zyh ; Lundhammar:2020xvw . In the present work, the spinless Salpeter equation (2) is not used to calculate the masses of bound states but rather to obtain the Regge trajectory relations. For simplicity, the diquark is regarded as pointlike.
Following Refs. Ferretti:2019zyh ; Bedolla:2019zwg ; Ferretti:2011zz ; Eichten:1974af , we employ the potential
$V_{d,t,p}$ $=-\frac{3}{4}\left[V_{c}+{\sigma}r+C\right]\left({\bf{F}_{i}}\cdot{\bf{F}_{j}}\right)_{d,t,p},$ (4)
where $V_{c}\propto{1/r}$ is a color Coulomb potential or a Coulomb-like potential arising from one-gluon-exchange. Here, $\sigma$ is the string tension, and $C$ is a fundamental parameter Gromes:1981cb ; Lucha:1991vn . The expression in brackets is the Cornell potential Eichten:1974af . The first two terms in Eq. (4) are determined solely by the nature of the color field; therefore, the same potential used for mesons and baryons Godfrey:1985xj ; Capstick:1986ter can also be used to describe the interaction between diquark and triquark Lebed:2015tna ; Brodsky:2014xia ; Song:2024bkj . ${\bf{F}_{i}}\cdot{\bf{F}_{j}}$ is the color-Casimir, with
$\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{d,t}\rangle=-\frac{2}{3},\quad\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{p}\rangle=-\frac{4}{3}.$ (5)
For heavy-heavy systems, where $m_{1},m_{2}{\gg}{{|\bf p|}}$, Eq. (2) reduces to
$M\Psi_{d,t,p}({{\bf r}})=\left[(m_{1}+m_{2})+\frac{{{\bf p}}^{2}}{2\mu}+V_{d,t,p}\right]\Psi_{d,t,p}({{\bf r}}),$ (6)
where
$\mu=m_{1}m_{2}/(m_{1}+m_{2}).$ (7)
By employing the Bohr-Sommerfeld quantization approach brau:04bs and using Eqs. (4) and (6), we obtain the following parametrized relation Chen:2022flh ; Chen:2021kfw
$M=$ $m_{R}+\beta_{x}(x+c_{0x})^{2/3},$
$x=$ $l_{1},l_{2},\,n_{r_{1}},\,n_{r_{2}},\,L_{1},\,L_{2},\,N_{r_{1}}\,N_{r_{2}},$ (8)
with
$\beta_{x}=c_{fx}c_{x}c_{c},\quad m_{R}=m_{1}+m_{2}+C^{\prime},$ (9)
where
$C^{\prime}=\left\{\begin{array}[]{cc}C/2,&\text{diquarks,\,triquarks},\\ C,&\text{pentaquarks}.\end{array}\right.$ (10)
$\sigma^{\prime}=\left\{\begin{array}[]{cc}\sigma/2,&\text{diquarks,\,triquarks},\\ \sigma,&\text{pentaquarks}.\end{array}\right.$ (11)
$c_{x}$ and $c_{c}$ are
$c_{c}=\left(\frac{\sigma^{\prime 2}}{\mu}\right)^{1/3},$ $\quad c_{l_{1},l_{2},L_{1},L_{2}}=\frac{3}{2},$
$c_{n_{r_{1}},n_{r_{2}},N_{r_{1}},N_{r_{2}}}$ $=\frac{\left(3\pi\right)^{2/3}}{2}.$ (12)
Theoretically, $c_{fx}$ equals one but int practice is fitted to data. In Eq. (II.2), $m_{1}$, $m_{2}$, $c_{x}$ and $\sigma$ are universal parameters for heavy-heavy systems. $c_{0x}$ varies with different Regge trajectories.
Using Eqs. (II.2), (9), (10), (11), and (II.2), we obtain the Regge trajectory relations for fully heavy pentaquarks:
$M$ $=m_{R_{\lambda_{1}}}+\beta_{x_{\lambda_{1}}}(x_{\lambda_{1}}+c_{0x_{\lambda_{1}}})^{2/3}\;(x_{\lambda_{1}}=L_{1},\,N_{r_{1}}),$
$m_{R_{\lambda_{1}}}$ $=M_{d_{1}}+M_{t}+C,$
$M_{d_{1}}$ $=m_{R_{\rho_{1}}}+\beta_{x_{\rho_{1}}}(x_{\rho_{1}}+c_{0x_{\rho_{1}}})^{2/3}\;(x_{\rho_{1}}=l_{1},\,n_{r_{1}}),$
$m_{R_{\rho_{1}}}$ $=m_{q_{1}}+m_{q_{2}}+C/2,$
$M_{t}$ $=m_{R_{\lambda_{2}}}+\beta_{x_{\lambda_{2}}}(x_{\lambda_{2}}+c_{0x_{\lambda_{2}}})^{2/3}\;(x_{\lambda_{2}}=L_{2},\,N_{r_{2}}),$
$m_{R_{\lambda_{2}}}$ $=M_{d_{2}}+m_{q_{3}}+C/2,$
$M_{d_{2}}$ $=m_{R_{\rho_{2}}}+\beta_{x_{\rho_{2}}}(x_{\rho_{2}}+c_{0x_{\rho_{2}}})^{2/3}\;(x_{\rho_{2}}=l_{2},\,n_{r_{2}}),$
$m_{R_{\rho_{2}}}$ $=m_{q_{4}}+m_{q_{5}}+C/2,$ (13)
where
$\mu_{\lambda_{1}}$ $=\frac{M_{d_{1}}M_{t}}{M_{d_{1}}+M_{t}},\;\mu_{\rho_{1}}=\frac{m_{q_{1}}m_{q_{2}}}{m_{q_{1}}+m_{q_{2}}},$
$\beta_{L_{1}}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda_{1}}}\right)^{1/3}c_{fL_{1}},\;\beta_{N_{r_{1}}}=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda_{1}}}\right)^{1/3}c_{fN_{r_{1}}},$
$\beta_{l_{1}}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{1}}}\right)^{1/3}c_{fl_{1}},\;\beta_{n_{r_{1}}}=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{1}}}\right)^{1/3}c_{fn_{r_{1}}},$
$\mu_{\lambda_{2}}$ $=\frac{M_{d_{2}}m_{q_{3}}}{M_{d_{2}}+m_{q_{3}}},\;\mu_{\rho_{2}}=\frac{m_{q_{4}}m_{q_{5}}}{m_{q_{4}}+m_{q_{5}}},$
$\beta_{L_{2}}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{4\mu_{\lambda_{2}}}\right)^{1/3}c_{fL_{2}},\;\beta_{N_{r_{2}}}=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{4\mu_{\lambda_{2}}}\right)^{1/3}c_{fN_{r_{2}}},$
$\beta_{l_{2}}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{2}}}\right)^{1/3}c_{fl_{2}},\;\beta_{n_{r_{2}}}=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{2}}}\right)^{1/3}c_{fn_{r_{2}}}.$ (14)
Here, $M$, $M_{t}$, $M_{d_{1}}$, $M_{d_{2}}$, $m_{q_{1}}$, $m_{q_{2}}$, $m_{q_{3}}$, $m_{q_{4}}$, and $m_{q_{5}}$ are the masses of the pentaquark, triquark, diquark $1$, diquark $2$, quark $1$, quark $2$, quark $3$, quark $4$, and quark $5$, respectively.
In Eq. (II.2), $M_{d_{1}}$ and $M_{d_{2}}$ give the diquark Regge trajectories Feng:2023txx , $M_{d_{1}}$, $M_{d_{2}}$ and $M_{t}$ can can be used to construct the triquark Regge trajectories Song:2024bkj . $M_{d_{1}}$, $M_{d_{2}}$, $M_{t}$ and $M$ generate the Regge trajectories for the fully heavy pentaquarks. According to Eqs. (II.2) and (II.2), we have
$M=M_{d_{1}}+M_{t}+C+\beta_{x_{\lambda_{1}}}(x_{\lambda_{1}}+c_{0x_{\lambda_{1}}})^{2/3}\;(x_{\lambda_{1}}=L_{1},\,N_{r_{1}})$ (15)
when diquark $1$ and the triquark are regarded as structureless constituents, the binding energies of the pentaquark is $\epsilon=C+\beta_{x_{\lambda_{1}}}(x_{\lambda_{1}}+c_{0x_{\lambda_{1}}})^{2/3}$. When the triquark is taken as a bound state composed a diquark and an antiquark, and diquarks $1$ and $2$ are each regarded as bound states of two heavy quarks, we have
$M=$ $m_{q_{1}}+m_{q_{2}}+m_{q_{3}}+m_{q_{4}}+m_{q_{5}}+\frac{5}{2}C$
$+\beta_{x_{\lambda_{1}}}(x_{\lambda_{1}}+c_{0x_{\lambda_{1}}})^{2/3}+\beta_{x_{\rho_{1}}}(x_{\rho_{1}}+c_{0x_{\rho_{1}}})^{2/3}$
$+\beta_{x_{\lambda_{2}}}(x_{\lambda_{2}}+c_{0x_{\lambda_{2}}})^{2/3}+\beta_{x_{\rho_{2}}}(x_{\rho_{2}}+c_{0x_{\rho_{2}}})^{2/3}.$ (16)
From the pentaquark Regge trajectory relations [Eq. (II.2)], it can be seen that there are four series of masses and, correspondingly, four series of Regge trajectories for fully heavy pentaquarks: the $\lambda_{1}$-trajectories (with $x_{\lambda_{2}}$, $x_{\rho_{1}}$, and $x_{\rho_{2}}$ fixed); the $\lambda_{2}$-trajectories (with $x_{\lambda_{1}}$, $x_{\rho_{1}}$, and $x_{\rho_{2}}$ fixed); the $\rho_{1}$-trajectories (with $x_{\lambda_{1}}$, $x_{\lambda_{2}}$, and $x_{\rho_{2}}$ fixed); and the $\rho_{2}$-trajectories (with $x_{\lambda_{1}}$, $x_{\lambda_{2}}$, and $x_{\rho_{1}}$ fixed).
For later convenience, we refer to the Regge trajectories calculated from Eqs. (II.2) and (II.2), or from Eqs. (II.2) and (II.2), as the complete forms of the Regge trajectories. The obtained constant and the mode under consideration are referred to the main part of the Regge trajectories. For example, when considering the $\rho_{1}$-trajectories, $\beta_{x_{\lambda_{2}}}(x_{\lambda_{2}}+c_{0x_{\lambda_{2}}})^{2/3}$ and $\beta_{x_{\rho_{2}}}(x_{\rho_{2}}+c_{0x_{\rho_{2}}})^{2/3}$ are constants, while $\beta_{x_{\lambda_{1}}}(x_{\lambda_{1}}+c_{0x_{\lambda_{1}}})^{2/3}$ becomes a function of $x_{\rho_{1}}$ (through the dependence in $\beta_{x_{\lambda_{1}}}$). Therefore, the main part of the $\rho_{1}$-trajectories is $\widetilde{m}_{R}+\beta_{x_{\rho_{1}}}(x_{\rho_{1}}+c_{0x_{\rho_{1}}})^{2/3}$, where $\widetilde{m}_{R}=m_{q_{1}}+m_{q_{2}}+m_{q_{3}}+m_{q_{4}}+m_{q_{5}}+\frac{5}{2}C+\beta_{x_{\lambda_{2}}}(x_{\lambda_{2}}+c_{0x_{\lambda_{2}}})^{2/3}+\beta_{x_{\rho_{2}}}(x_{\rho_{2}}+c_{0x_{\rho_{2}}})^{2/3}$. The difference between the complete form of the $\rho_{1}$-trajectory and its main part is $\beta_{x_{\lambda_{1}}}(x_{\lambda_{1}}+c_{0x_{\lambda_{1}}})^{2/3}$.
III Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$
III.1 Parameters
Table 1: The values of parameters Feng:2023txx ; Faustov:2021qqf .
$m_{c}=1.55\;{{\rm GeV}}$,   $m_{b}=4.88\;{{\rm GeV}}$,
$\sigma=0.18\;{{\rm GeV}^{2}}$,  $C=-0.3\;{{\rm GeV}}$,
$c_{fn_{r}}=1.0$,  $c_{fl}=1.17$
$(cc)$ $c_{0n_{r}}(1^{3}s_{1})=0.205$, $c_{0{l}}(1^{3}s_{1})=0.337$,
$(bb)$ $c_{0n_{r}}(1^{3}s_{1})=0.01$, $c_{0{l}}(1^{3}s_{1})=0.001$.
The parameter values are listed in Table 1. The values of $m_{b}$, $m_{c}$, $\sigma$ and $C$ are adopted directly from Ref. Faustov:2021qqf . $c_{fx}$ and $c_{0x}$ for the $\rho$-mode are extracted by fitting the Regge trajectories for doubly heavy mesons, and are subsequently used as input to fit the Regge trajectories for doubly heavy diquarks. The parameter $c_{fx}$ is universal for all doubly heavy diquark Regge trajectories, whereas $c_{0x}$ varies with different diquark Regge trajectories Feng:2023txx . The parameters for the $\lambda$-mode are determined by the following relations Xie:2024dfe
$c_{fL_{1},fL_{2}}=$ $1.116+0.013\mu_{\lambda_{1},\lambda_{2}},$
$c_{0L_{1},0L_{2}}=$ $0.540-0.141\mu_{\lambda_{1},\lambda_{2}},$
$c_{fN_{r_{1}},fN_{r_{2}}}=$ $1.008+0.008\mu_{\lambda_{1},\lambda_{2}},$
$c_{0N_{r_{1}},0N_{r_{2}}}=$ $0.334-0.087\mu_{\lambda_{1},\lambda_{2}},$ (17)
where $\mu_{\lambda_{1}}$ and $\mu_{\lambda_{2}}$ are the reduced masses (see Eq. (II.2)). The relations in Eq. (17) are obtained by fitting the mesons, baryons, and tetraquarks. These relations are employed as a provisional method until a better approach becomes available. Validation can be achieved by comparing the fitted results for pentaquarks with theoretical values obtained from other approaches and available experimental data.
III.2 $\rho$-trajectories for the pentaquark $P_{cc\bar{c}bb}$
Table 2: The masses of the radially and orbitally $\rho_{1}$- and $\rho_{2}$-excited states of the pentaquark $P_{cc\bar{c}bb}$ (in ${{\rm GeV}}$). Both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ are considered. The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}^{2s_{3}+1}L_{2J_{2}},N_{1}^{2s_{4}+1}L_{1J_{1}}\rangle$. $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ denotes the spin-averaged states. Eqs. (II.2), (II.2) and (17) are used in the calculation. States marked with (×)($\times$)( × ) do not exist.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ $(cc)(\bar{c}(bb))$ $(bb)(\bar{c}(cc))$
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|2^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.49 14.41
$|3^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.75 14.61
$|4^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.98 14.77
$|5^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 15.18 14.91
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|1^{3}s_{1},2^{3}s_{1},1S,1S\rangle$ 14.45 14.46
$|1^{3}s_{1},3^{3}s_{1},1S,1S\rangle$ 14.64 14.72
$|1^{3}s_{1},4^{3}s_{1},1S,1S\rangle$ 14.80 14.95
$|1^{3}s_{1},5^{3}s_{1},1S,1S\rangle$ 14.95 15.15
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.14 14.08
$|1^{3}p_{2},1^{3}s_{1},1S,1S\rangle(\times)$ 14.41 14.34
$|1^{3}d_{3},1^{3}s_{1},1S,1S\rangle$ 14.61 14.49
$|1^{3}f_{4},1^{3}s_{1},1S,1S\rangle(\times)$ 14.78 14.62
$|1^{3}g_{5},1^{3}s_{1},1S,1S\rangle$ 14.94 14.74
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.12 14.11
$|1^{3}s_{1},1^{3}p_{2},1S,1S\rangle(\times)$ 14.37 14.38
$|1^{3}s_{1},1^{3}d_{3},1S,1S\rangle$ 14.53 14.58
$|1^{3}s_{1},1^{3}f_{4},1S,1S\rangle(\times)$ 14.66 14.75
$|1^{3}s_{1},1^{3}g_{5},1S,1S\rangle$ 14.77 14.91
Figure 2: The $\rho_{1}$- and $\rho_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$. $n_{r_{1}}$ and ${l_{1}}$ are the radial and orbital quantum numbers for the $\rho_{1}$-mode, respectively, while $n_{r_{2}}$ and ${l_{2}}$ are the corresponding numbers for the $\rho_{2}$-mode. Circles represent the predicted data listed in Table 2. The black dashed lines correspond to the $\rho$-trajectories for the complete forms, obtained from Eqs. (II.2) and (II.2) or from Eqs. (II.2) and (II.2). The pink lines correspond to the fitted formulas, obtained by linearly fitting the predicted data in Table 2; these formulas are listed in Table 5. The blue dotted lines correspond to the main parts of the complete form, which are also listed in Table 5.
In this subsection, the pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ is considered. By using Eqs. (II.2), (II.2) and (17) and parameters in Table 1, the spin-averaged masses of the radially and orbitally $\rho_{1}$- and $\rho_{2}$-excited states are calculated. The calculated results are listed in Table 2. More results can be easily obtained in a similar manner. When calculating the $\rho_{1}$-modes, all other modes are kept in their ground states, and the parameters used correspond to the radial ground states. By a similar procedure, the spin-averaged masses of the $\rho_{2}$-excited states can be determined.
In the configuration $(cc)(\bar{c}(bb))$, the masses of the radially and orbitally $\rho_{1}$-excited states are greater than those of the corresponding $\rho_{2}$-excited states (see Table 2). Conversely, in the configuration $(bb)(\bar{c}(cc))$, the masses of the $\rho_{2}$-excited states exceed those of the $\rho_{1}$-excited states. This is because the masse increase from excitation of the diquark $(cc)$ is greater than that from excitation of the diquark $(bb)$. Accordingly, the following inequalities hold for the pentaquark $P_{cc\bar{c}bb}$:
$M\left(\rho_{1},(cc)(\bar{c}(bb))\right)>M\left(\rho_{2},(cc)(\bar{c}(bb))\right),$
$M\left(\rho_{2},(bb)(\bar{c}(cc))\right)>M\left(\rho_{1},(bb)(\bar{c}(cc))\right),$
$M\left(\rho_{1},(cc)(\bar{c}(bb))\right)>M\left(\rho_{1},(bb)(\bar{c}(cc))\right),$
$M\left(\rho_{2},(bb)(\bar{c}(cc))\right)>M\left(\rho_{2},(cc)(\bar{c}(bb))\right),$ (18)
where $M\left(\rho_{1},(bb)(\bar{c}(cc))\right)$ denotes the masses of the radially or orbitally $\rho_{1}$-mode excited state of $P_{cc\bar{c}bb}$ in the configuration $(bb)(\bar{c}(cc))$; the other notations in Eq. (III.2) are similarly defined. The masses of the $\rho_{1}$-excited ($\rho_{2}$-excited) states in the configuration $(cc)(\bar{c}(bb))$ are close to those of the $\rho_{2}$-excited ($\rho_{1}$-excited) states in the configuration $(bb)(\bar{c}(cc))$, i.e., $M\big{(}\rho_{1},(cc)(\bar{c}(bb))\big{)}{\approx}M\big{(}\rho_{2},(bb)(\bar{c}(cc))\big{)}$, $M\big{(}\rho_{1},(bb)(\bar{c}(cc))\big{)}{\approx}M\big{(}\rho_{2},(cc)(\bar{c}(bb))\big{)}$. Possible mixings between these states are neglected here.
Inspired by Ref. Burns:2010qq , we plot the $\rho_{1}$- and $\rho_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ in the $(M,\,(x+c_{0})^{2/3})$ plane (see Fig. 2). The full forms of the Regge trajectories are directly calculated from Eqs. (II.2) and (II.2), or from Eqs. (II.2) and (II.2). As examples, the full forms of the $\rho_{1}$-trajectories and $\rho_{2}$-trajectories are listed in Eqs. (A) and (A). These expressions are generally lengthy and cumbersome. By linearly fitting the calculated data in Table 2, we obtain the fitted formulas listed in Table 5. As shown in Fig. 2, the fitted formulas overlap with the complete forms of the $\rho_{1}$- and $\rho_{2}$-trajectories. This indicates that the complicated full forms of the $\rho$-Regge trajectories can be well approximated by the simple fitted formulas. Therefore, we have that the complete forms of the $\rho_{1}$- and $\rho_{2}$-trajectories all exhibit the behavior $M{\sim}x_{\rho_{1}}^{2/3},\;x_{\rho_{2}}^{2/3}$, where $x_{\rho_{1}}=n_{r_{1}},\;l_{1}$ and $x_{\rho_{2}}=n_{r_{2}},\;l_{2}$. The main parts of the full forms of the $\rho_{1}$- and $\rho_{2}$-trajectories [listed in Table 5] show significant deviations from the full forms. For both the radial and orbital Regge trajectories, and for both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$, the main parts of the $\rho_{2}$-trajectories deviate more from the complete forms than those of the $\rho_{1}$-trajectories.
III.3 $\lambda$-trajectories for the pentaquark $P_{cc\bar{c}bb}$
Table 3: Same as Table 2 except for the $\lambda_{1}$- and $\lambda_{2}$-excited states.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ $(cc)(\bar{c}(bb))$ $(bb)(\bar{c}(cc))$
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|1^{3}s_{1},1^{3}s_{1},2S,1S\rangle$ 14.45 14.43
$|1^{3}s_{1},1^{3}s_{1},3S,1S\rangle$ 14.68 14.67
$|1^{3}s_{1},1^{3}s_{1},4S,1S\rangle$ 14.88 14.88
$|1^{3}s_{1},1^{3}s_{1},5S,1S\rangle$ 15.06 15.06
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|1^{3}s_{1},1^{3}s_{1},1S,2S\rangle$ 14.58 14.54
$|1^{3}s_{1},1^{3}s_{1},1S,3S\rangle$ 14.89 14.83
$|1^{3}s_{1},1^{3}s_{1},1S,4S\rangle$ 15.15 15.08
$|1^{3}s_{1},1^{3}s_{1},1S,5S\rangle$ 15.39 15.30
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|1^{3}s_{1},1^{3}s_{1},1P,1S\rangle$ 14.36 14.33
$|1^{3}s_{1},1^{3}s_{1},1D,1S\rangle$ 14.53 14.51
$|1^{3}s_{1},1^{3}s_{1},1F,1S\rangle$ 14.67 14.66
$|1^{3}s_{1},1^{3}s_{1},1G,1S\rangle$ 14.80 14.80
$|1^{3}s_{1},1^{3}s_{1},1S,1S\rangle$ 14.13 14.10
$|1^{3}s_{1},1^{3}s_{1},1S,1P\rangle$ 14.45 14.42
$|1^{3}s_{1},1^{3}s_{1},1S,1D\rangle$ 14.68 14.64
$|1^{3}s_{1},1^{3}s_{1},1S,1F\rangle$ 14.88 14.82
$|1^{3}s_{1},1^{3}s_{1},1S,1G\rangle$ 15.06 14.99
Figure 3: The $\lambda_{1}$- and $\lambda_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$. $N_{r_{1}}$ and ${L_{1}}$ are the radial and orbital quantum numbers for the $\lambda_{1}$-mode, respectively. $N_{r_{2}}$ and ${L_{2}}$ are the radial and orbital quantum numbers for the $\lambda_{2}$-mode, respectively. Circles represent the predicted data listed in Table 3. The black dashed lines correspond to the $\lambda$-trajectories for the complete forms, obtained from Eqs. (II.2) and (II.2) or from Eqs. (II.2) and (II.2). The pink lines correspond to the fitted formulas, obtained by linearly fitting the calculated data in Table 3; these formulas are listed in Table 5. The blue dotted lines are for the main parts of the full forms, which are also listed in Table 5.
In this subsection, the $\lambda_{1}$-mode and $\lambda_{2}$-mode of the pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ are considered. Using Eqs. (II.2), (II.2) and (17), together with the parameters in Table 1, the masses of the radially and orbitally $\lambda_{1}$- and $\lambda_{2}$-excited states are calculated. The calculated results are listed in Table 3. When calculating the $\lambda_{1}$ modes, all other modes are kept in their ground states, and the parameters used correspond to the radial ground states of those modes. By a similar procedure, the masses of the $\lambda_{2}$-excited states can be determined.
In both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$, the masses of the radially and orbitally $\lambda_{1}$-excited states are greater than those of the corresponding $\lambda_{2}$-excited states (see Table 3). Accordingly, the following relations hold for the pentaquark $P_{cc\bar{c}bb}$
$M\left(\lambda_{1},(cc)(\bar{c}(bb))\right)>M\left(\lambda_{2},(cc)(\bar{c}(bb))\right),$
$M\left(\lambda_{1},(cc)(\bar{c}(bb))\right)>M\left(\lambda_{2},(bb)(\bar{c}(cc))\right),$
$M\left(\lambda_{1},(bb)(\bar{c}(cc))\right)>M\left(\lambda_{2},(bb)(\bar{c}(cc))\right),$
$M\left(\lambda_{1},(bb)(\bar{c}(cc))\right)>M\left(\lambda_{2},(cc)(\bar{c}(bb))\right),$ (19)
where $M\left(\lambda_{1},(bb)(\bar{c}(cc))\right)$ denotes the masses of the radially or orbitally $\lambda_{1}$-excited state of $P_{cc\bar{c}bb}$ in configuration $(bb)(\bar{c}(cc))$; the other notations in Eq. (III.3) are similarly defined. The masses of the $\lambda_{1}$-excited ($\lambda_{2}$-excited) states in the configuration $(cc)(\bar{c}(bb))$ are close to those of the $\lambda_{1}$-excited ($\lambda_{2}$-excited) states in the configuration $(bb)(\bar{c}(cc))$, i.e., $M\big{(}\lambda_{1},(cc)(\bar{c}(bb))\big{)}{\approx}M\big{(}\lambda_{1},(bb)(\bar{c}(cc))\big{)}$, $M\big{(}\lambda_{2},(cc)(\bar{c}(bb))\big{)}{\approx}M\big{(}\lambda_{2},(bb)(\bar{c}(cc))\big{)}$. Possible mixings among these states are not considered here.
The $\lambda_{1}$- and $\lambda_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ are shown in Fig. 3. The complete forms of the $\lambda$-Regge trajectories can be calculated easily from Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2). Since the diquark $1$ and the triquark are treated as individual constituents without considering their internal substructures, the complete form of the $\lambda_{1}$-trajectories coincides with its main part. Thus, the $\lambda_{1}$-trajectories are the simplest among the four series of Regge trajectories (see Table 5 and Fig. 3). In Fig. 3, the graphs of the compete forms, the fitted formulas, and the main parts of the $\lambda_{1}$-trajectories overlap. Similar to the $\rho$-trajectories, the explicit forms of the $\lambda_{2}$-trajectories are lengthy. By performing a linear fit to the calculated data, we obtain the fitted formulas presented in Table 5. As shown in Fig. 3, these fitted formulas closely overlap with the complete forms of the $\lambda_{2}$-trajectories, indicating that the full forms can be well approximated by the simple fitted formulas. Therefore, both the complete forms and the fitted formulas for the $\lambda_{2}$-trajectories exhibit the behavior $M{\sim{x_{\lambda_{2}}^{2/3}}}$. However, the main parts of the complete forms of the $\lambda_{2}$-trajectories [listed in Table 5] deviate from the full expressions.
Table 4: Comparison of theoretical predictions for the spin-averaged masse of the ground state of pentaquark $P_{cc\bar{c}bb}$ (in GeV).
Mass
$\{cc\}(\bar{c}\{bb\})$ 14.13
$\{bb\}(\bar{c}\{cc\})$ 14.10
$\{bc\}(\bar{c}\{bc\})$ 14.13
Ref. Rashmi:2024ako 15.21
Ref. Liang:2024met 14.62
Ref. Gordillo:2024blx 14.30
Ref. An:2022fvs 14.57
Ref. Zhang:2023hmg 14.86
IV Discussions
By employing the Regge trajectory formulas in Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2), the masses of higher excited states can be easily calculated. We argue that calculating the masses of highly excited states is instructive, even though these states are expected to be unstable. This is because various models or approaches for calculating the masses of pentaquarks often employ different parameter values. Nevertheless, they generally provide similar predictions for the ground state and lower excited states. However, the theoretical predictions for the higher excited states are expected to differ more significantly, which can be instructive.
One merit of the Regge trajectory approach is its simple analytical form. However, the complete forms of the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories for pentaquarks are quite lengthy and cumbersome (see Eqs. (A) and (A)). We fit the calculated data to obtain simplified relations, as listed in Table 5. Because the used data are calculated by using the complete forms of the Regge trajectories, the lengthy complete forms can be well approximated by the simple fitted relations, and they have the same behaviors (see Figs. 2 and 3).
It is interesting that the main parts of the complete forms and the fitted formulas have the same functional form $M=m_{R}+\beta(x+c_{0})^{2/3}$. However, the values of the parameters $m_{R}$ and $\beta$ change significantly. This means that, unlike in the meson case, an obvious and direct dependence of $m_{R}$ and $\beta$ on the constituents’ masses and string tension is no longer present when we construct simple $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectory formulas for pentaquarks. It can be concluded from the preceding discussions, as well as those in Sec. III, that the $\rho_{1}$, $\rho_{2}$, and $\lambda_{2}$ Regge trajectories for fully heavy pentaquarks cannot be obtained merely by mimicking the meson Regge trajectories; instead, they should be constructed based on the pentaquark’s structure and substructure. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will remain unclear. The predictive power of the Regge trajectories would be compromised.
The Regge trajectories assume different forms and exhibit distinct behaviors across different energy regions Chen:2022flh ; Chen:2021kfw . The diquark $(bb)$, $(cc)$ and the triquark $(\bar{c}(cc))$, $(\bar{c}(bb))$ are all the heavy-heavy systems; therefore, the main parts of the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-trajectories all follow the behavior $M{\sim}x^{2/3}$, where $x=x_{\rho_{1}},\,x_{\rho_{2}},\,x_{\lambda_{1}},\,x_{\lambda_{2}}$ (see Eq. (II.2)). Fitting the calculated data shows that the complete forms of the Regge trajectories also exhibit the same behavior as their main parts, namely $M{\sim}x^{2/3}$, see Table 5.
For fully heavy pentaquarks, all four series of Regge trajectories are concave downward in the $(M,\,x)$ planes when the confining potential is linear. It can also be easily shown that these trajectories are concave downward in the $(M^{2},\,x)$ planes when the confining potential is linear.
To our knowledge, this is the first systematic discussion of all four series of Regge trajectories for fully heavy pentaquarks (see also Ref. Song:2024bkj , which cites the results of this work). Previously, the relation in Eq. (II.2), the potential in Eq. (4), and the parameter values in Table 1 have been successfully applied to discuss the Regge trajectories for mesons, baryons, diquarks, triquarks, and tetraquarks Chen:2023djq ; Chen:2022flh ; Feng:2023txx ; Xie:2024lfo ; Xie:2024dfe ; Song:2024bkj . They are now employed to study pentaquark $P_{cc\bar{c}bb}$, yielding results consistent with other theoretical predictions (see Table 4). This not only demonstrates the universality of the Regge trajectory relation and parameter values, but also illustrates its predictive capability.
V Conclusions
In this work, we propose Regge trajectory relations for the fully heavy pentaquark $P_{cc\bar{c}bb}$, employing both diquark and triquark Regge trajectories. By employing these new relations, we present four series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$: the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-trajectories. The masses of the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-excited states are roughly estimated.
The complete forms of the Regge trajectories for pentaquark $P_{cc\bar{c}bb}$ are often lengthy and cumbersome. Except for the $\lambda_{1}$-trajectories, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories for the fully heavy pentaquarks cannot be obtained by simply mimicking the meson Regge trajectories but should instead be constructed based on pentaquark’s structure and substructure. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will remain unclear. The predictive power of the Regge trajectories would be compromised.
We show that the lengthy complete forms of the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can be well approximated by the simple fitted formulas. All four series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ exhibit a behavior of $M{\sim}x^{2/3}$, where $x=n_{r_{1}},n_{r_{2}},l_{1},l_{2},N_{r_{1}},N_{r_{2}},L_{1},L_{2}$. Moreover, all four series of trajectories are concave downward in the $(M^{2},\,x)$ plane.
Appendix A List of the Regge trajectory relations
The concrete forms of the Regge trajectories for the pentaquark are not as same simple as those for mesons. Although Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2) are simple, their final forms can be rather tedious, because pentaquarks have substructures and the slope depends on the masses of the constituents.
From Eqs. (II.2) and (II.2) or Eqs. (II.2) and (II.2), we can easily obtain the complete forms of the Regge trajectories for the fully heavy pentaquark $P_{cc\bar{c}bb}$. The resulting expressions are rather long and tedious. As an example, we list only the radial $\rho_{1}$- and $\rho_{2}$-trajectories for the configuration $(cc)(\bar{c}(bb))$. The $\rho_{1}$-trajectory for the configuration $(cc)(\bar{c}(bb))$ is given by
$M=$ $13.8255+0.487766(0.205+n_{r_{1}})^{2/3}+0.318109\left(\frac{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}{2.95+0.487766(0.205+n_{r_{1}})^{2/3}}\right)^{1/3}$
×$\times\left(0.334-\frac{0.97227\left(2.95+0.487766(0.205+n_{r_{1}})^{2/3}\right)}{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}\right)^{2/3}\left(1.008+\frac{0.0894041\left(2.95+0.487766(0.205+n_{r_{1}})^{2/3}\right)}{14.1255+0.487766(0.205+n_{r_{1}})^{2/3}}\right).$ (20)
The $\rho_{2}$-trajectory for the configuration $(cc)(\bar{c}(bb))$ reads
$M=$ $13.8296+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\left(\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\right)^{1/3}\left(0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\right)^{2/3}$
×$\times\left(1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\right)+0.486744\Big{(}\Big{(}14.1296+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\Big{(}\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{1/3}\Big{(}0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{2/3}$
$\Big{(}1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}\Big{)}\Big{/}\Big{(}11.01+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\Big{(}\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{1/3}\Big{(}0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{2/3}$
$\Big{(}1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}\Big{)}\Big{)}^{1/3}\Big{(}0.334-\Big{(}0.271404\Big{(}11.01+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\Big{(}\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{1/3}\Big{(}0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{2/3}$
$\Big{(}1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}\Big{)}\Big{)}\Big{/}\Big{(}14.1296+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\Big{(}\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{1/3}\Big{(}0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{2/3}$
$\Big{(}1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}\Big{)}\Big{)}^{2/3}\Big{(}1.008+\Big{(}0.0249567\Big{(}11.01+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\Big{(}\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{1/3}\Big{(}0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{2/3}$
$\Big{(}1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}\Big{)}\Big{)}\Big{/}\Big{(}14.1296+0.332799(0.01+n_{r_{2}})^{2/3}$
$+0.38714\Big{(}\frac{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}{9.61+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{1/3}\Big{(}0.334-\frac{0.13485(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}^{2/3}$
$\Big{(}1.008+\frac{0.0124(9.61+0.332799(0.01+n_{r_{2}})^{2/3})}{11.16+0.332799(0.01+n_{r_{2}})^{2/3}}\Big{)}\Big{)}\Big{)}.$ (21)
The explicit forms of the Regge trajectories calculated from Eqs. (II.2) and (II.2) or from Eqs. (II.2) and (II.2) are often rather tedious. Here, we present only the main parts of the complete forms of the Regge trajectories, as well as the fitted formulas obtained by fitting the calculated results (see Table 5).
Table 5: The fitted formulas and the main parts of the $\rho$- and $\lambda$-Regge trajectories for pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}bb)$ and $(bb)(\bar{c}cc)$. $({\rm fit})$ denotes the formula obtained by fitting the calculated results. $({\rm main})$ indicates the main part of the complete forms of the Regge trajectories, derived from Eqs. (II.2) and (II.2).
$(cc)(\bar{c}(bb))$ $(bb)(\bar{c}(cc))$
$\rho_{1}$ $M=13.9661+0.465794(0.205+n_{r_{1}})^{2/3}$ (fit) $M=14.0817+0.329997(0.01+n_{r_{1}})^{2/3}$ (fit)
$M=13.8255+0.487766(0.205+n_{r_{1}})^{2/3}$ (main) $M=14.0053+0.332799(0.01+n_{r_{1}})^{2/3}$ (main)
$M=13.9663+0.366186(0.337+l_{1})^{2/3}$ (fit) $M=14.0817+0.259577(0.001+l_{1})^{2/3}$ (fit)
$M=13.8255+0.383712(0.337+l_{1})^{2/3}$ (main) $M=14.0053+0.261804(0.001+l_{1})^{2/3}$ (main)
$\rho_{2}$ $M=14.1131+0.331141(0.01+n_{r_{2}})^{2/3}$ (fit) $M=13.9352+0.465439(0.205+n_{r_{2}})^{2/3}$ (fit)
$M=13.8296+0.332799(0.01+n_{r_{2}})^{2/3}$ (main) $M=13.6754+0.487766(0.205+n_{r_{2}})^{2/3}$ (main)
$M=14.1131+0.26048(0.001+l_{2})^{2/3}$ (fit) $M=13.9351+0.366203(0.337+l_{2})^{2/3}$ (fit)
$M=13.8296+0.261804(0.001+l_{2})^{2/3}$ (main) $M=13.6754+0.383712(0.337+l_{2})^{(}2/3)$ (main)
$\lambda_{1}$ $M=13.9951+0.542908(0.121824+N_{r_{1}})^{2/3}$ (fit) $M=14.0207+0.500992(0.0594404+N_{r_{1}})^{2/3}$ (fit)
$M=13.9951+0.542908(0.121824+N_{r_{1}})^{2/3}$ (main) $M=14.0207+0.500992(0.0594404+N_{r_{1}})^{2/3}$ (main)
$M=13.9951+0.407734(0.196129+L_{1})^{2/3}$ (fit) $M=14.0207+0.377205(0.0950241+L_{1})^{2/3}$ (fit)
$M=13.9951+0.407734(0.196129+L_{1})^{2/3}$ (main) $M=14.0207+0.377205(0.0950241+L_{1})^{2/3}$ (main)
$\lambda_{2}$ $M=13.979+0.413004(0.217853+N_{r_{2}})^{2/3}$ (fit) $M=13.9281+0.433149(0.243911+N_{r_{2}})^{2/3}$ (fit)
$M=13.845+0.414497(0.217853+N_{r_{2}})^{2/3}$ (main) $M=13.845+0.450067(0.243911+N_{r_{2}})^{2/3}$ (main)
$M=13.979+0.308932(0.351762+L_{2})^{2/3}$ (fit) $M=13.9279+0.32383(0.393994+L_{2})^{2/3}$ (fit)
$M=13.845+0.310068(0.351762+L_{2})^{2/3}$ (main) $M=13.845+0.336311(0.393994+L_{2})^{2/3}$ (main)